OpenStudy (dopeeaf):

Help .? Medal & fan .! Segment RT has the endpoints (-3,4) and (5,2). What is its midpoint? (1,1) (2,2) (1,3) (4,3)

OpenStudy (priyar):

midpoint=(x,y) let the segment's end points be (x1,y1) and (x2,y2) then x=(x1 + x2)/2 y=(y1+y2)/2

OpenStudy (dopeeaf):

Ohh I see, so it would be (2,2)?

OpenStudy (priyar):

i think u forgot to divide by 2..

OpenStudy (dopeeaf):

Oh okay nvrm, (4,3)?

OpenStudy (priyar):

no..can u tell me how u got this ..